# 1]Independent Researcher

### v1

@@ 3.2 Data-driven parameter discovery to (2+ 1)-dimensional VC-CNLSE
    $+\frac{1}{N_{r}}\sum_{j=1}^{N_{r}}w_{r}^{(k)}\Big{(}\Big{\lvert}\mathcal{F}_{\hat{u}_{1,\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2}}}(\mathbb{X}^{j},t^{j})\Big{\rvert}^{2}+\Big{\lvert}\mathcal{F}_{\hat{v}_{1,\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2}}}(\mathbb{X}^{j},t^{j})\Big{\rvert}^{2}$
    $+\Big{\lvert}\mathcal{G}_{\hat{u}_{2,\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2}}}(\mathbb{X}^{j},t^{j})\Big{\rvert}^{2}+\Big{\lvert}\mathcal{G}_{\hat{v}_{2,\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2}}}(\mathbb{X}^{j},t^{j})\Big{\rvert}^{2}\Big{)}$
    $+Loss_{a_{1}}+Loss_{a_{2}},$
@@
    where $\boldsymbol{\theta}_{1}$ and $\boldsymbol{\theta}_{2}$ are the sets of trainable parameters for the first network, which outputs the real and imaginary parts of the decomposed solutions $\psi_{1}$ and $\psi_{2}$, and the second network, which outputs the predictions of the variable coefficients, respectively. $Loss_{a_{1}}$ and $Loss_{a_{2}}$ denote the slope recovery loss terms for the first and second networks, respectively. After $10,000$ iterations of the Adam optimizer and some iterations of the L-BFGS optimizer with clean data (noise = $0\%$), the $\mathbb{L}^{2}$ relative error of the variable coefficients $\alpha(t)$ and $\beta(t)$ decreased from $0.40\%$ to $0.28\%$ and $0.65\%$ to $0.45\%$, respectively. When we add noise levels of $5\%$, $10\%$, and $15\%$ into the training observations, the $\mathbb{L}^{2}$ relative error of the dispersion coefficient $\alpha(t)$ changes from $0.53\%$, $0.55\%$, and $0.57\%$ to $0.33\%$, $0.27\%$, and $0.33\%$, respectively, when we first use iterations of the Adam optimizer and then continue the optimization process with the L-BFGS optimizer. This scenario holds for the nonlinearity coefficient $\beta(t)$, where the $\mathbb{L}^{2}$ relative error changes from $0.61\%$, $0.67\%$, and $0.72\%$ to $0.42\%$, $0.40\%$, and $0.60\%$, when we add noise levels of $5\%$, $10\%$, and $15\%$ into the training observations, respectively. The predictions of the dispersion and nonlinearity coefficients $\alpha(t)$ and $\beta(t)$ under different noise levels are displayed in the left and middle panels of Fig. 16. One can see that our enhanced trained PINN model can accurately predict the dynamic behavior of both variable coefficients, even in the presence of noise intensities. The $\mathbb{L}^{2}$ relative errors of both the dispersion and nonlinearity variable coefficients $\alpha(t)$ and $\beta(t)$ for different noise intensities are illustrated in the right panel of Fig. 16, which shows that increasing the noise levels does not significantly impact prediction accuracy. The obtained results show that our enhanced PINN model can accurately predict the behavior of the variable coefficients $\alpha(t)$ and $\beta(t)$ in the system of model equations (1) with a small amount of training data. It should be noted that these coefficients are nonlinear functions of the temporal variable. These results confirm that the proposed algorithm is robust even when noise levels are added to the training observations. In summary, the proposed enhanced PINN framework introduces a promising modeling approach in the field of optical fiber communication by accurately capturing the dynamic behavior of variable coefficients in nonlinear systems. Its robustness against noisy data makes it a reliable tool for real world applications, where measurement uncertainties are common. The ability to maintain high prediction accuracy under different noise intensities highlights the potential of this framework for parameter discovery and system identification in complex, data-driven environments.
@@
    Figure 16: The predictions of the dispersion coefficient $\alpha(t)$ (left panel) and the nonlinearity coefficient $\beta(t)$ (middle pannel) and the related prediction errors (right panel) under different noise intensities.
@@ 4 Conclusions
In this study, we proposed an enhanced PINN framework for data-driven soliton solutions and parameter discovery to the (2+ 1)-dimensional VC-CNLSE. We introduced a locally adaptive activation function to our PINN algorithm to accelerate the convergence speed of the training process. PINN algorithms enhance the capabilities of neural networks by incorporating the physical knowledge, described by differential equations, into the training phase. Incorporating physical knowledge into the training phase allows the network to learn from both data and the underlying principles governing the problem, leading to more accurate and physically consistent predictions. By leveraging the residual-based adaptive refinement (RAR) strategy, we introduced a region-specific weighted loss function, demonstrating that different distributions of physical knowledge are required when training a PINN model for the (2+ 1)-dimensional VC-CNLSE. We demonstrated the framework’s effectiveness in capturing the dynamics of vector dark and anti-dark one- and two-soliton structures through various soliton solutions and the corresponding data-driven predictions. The density plots of the exact and predicted solutions are displayed, and the related evolution dynamics are illustrated through three-dimensional plots. We also compared the exact and predicted solutions at different temporal levels to show the accuracy of the trained model’s prediction. We performed data-driven parameter discovery for the (2+ 1)-dimensional VC-CNLSE, categorizing it into two classes: constant parameter discovery and variable parameter discovery. For constant parameter discovery, we employed an enhanced PINN model equipped with a single neural network to estimate the value of the constant parameter in the system of model equations. For variable parameter discovery, we used a dual-network strategy within the proposed PINN framework to capture the dynamic behavior of the variable dispersion and nonlinearity coefficients. As these coefficients are nonlinear functions of the temporal variable, this shows that our enhanced PINN framework is capable of capturing complex nonlinear variable coefficients. This robustness highlights the framework’s potential for broader applications in complex optical fiber systems. The obtained results confirm that the proposed enhanced PINN framework offers a powerful and versatile approach for solving high-dimensional, complex solitonic dynamics in optical communication systems. Its robustness, accuracy, and adaptability make it a promising tool for advancing research and applications in nonlinear optics and related fields.

### v2

@@ 3.3 Data-driven parameter discovery to (2+ 1)-dimensional VC-CNLSE
$+\frac{1}{N_{r}}\sum_{k=1}^{N_{r}}\Big{(}w_{r}^{\mathcal{F}_{\hat{u}}}\Big{\lvert}\mathcal{F}_{\hat{u}_{1,\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2}}}(\mathbb{x}^{k},t^{k})\Big{\rvert}^{2}+w_{r}^{\mathcal{F}_{\hat{v}}}\Big{\lvert}\mathcal{F}_{\hat{v}_{1,\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2}}}(\mathbb{x}^{k},t^{k})\Big{\rvert}^{2}$
$+w_{r}^{\mathcal{G}_{\hat{u}}}\Big{\lvert}\mathcal{G}_{\hat{u}_{2,\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2}}}(\mathbb{x}^{k},t^{k})\Big{\rvert}^{2}+w_{r}^{\mathcal{G}_{\hat{v}}}\Big{\lvert}\mathcal{G}_{\hat{v}_{2,\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2}}}(\mathbb{x}^{k},t^{k})\Big{\rvert}^{2}\Big{)}$
$+\frac{1}{N_{vc}}\sum_{k=1}^{N_{vc}}{\Big{(}}{\Big{\lvert}}\alpha_{\boldsymbol{\theta}_{2}}(t_{k})-\alpha(t_{k}){\Big{\rvert}}^{2}+{\Big{\lvert}}\beta_{\boldsymbol{\theta}_{2}}(t_{k})-\beta(t_{k}){\Big{\rvert}}^{2}{\Big{)}}$
$+w_{a_{1}}\cdot Loss_{a_{1}}(\boldsymbol{\theta}_{1})+w_{a_{2}}\cdot Loss_{a_{2}}(\boldsymbol{\theta}_{2}),$
@@
where $\boldsymbol{\theta}_{1}$ and $\boldsymbol{\theta}_{2}$ are the sets of trainable parameters for the first network, which outputs the real and imaginary parts of the decomposed solutions $\psi_{1}$ and $\psi_{2}$, and the second network, which outputs the predictions of the variable coefficients, respectively. $Loss_{a_{1}}$ and $Loss_{a_{2}}$ denote the slope recovery loss terms for the first and second networks, respectively, and $w_{a_{1}}$ and $w_{a_{2}}$ are the corresponding weight coefficients. For this example, we set $w_{a_{1}}=1e-02$ and $w_{a_{2}}=1e-01$. The set $\Big{\{}t_{k},\alpha(t_{k}),\beta(t_{k})\Big{\}}_{k=1}^{N_{vc}}$ denotes the set of training data related to the boundary of the temporal interval. After $10,000$ iterations of the Adam optimizer and some iterations of the L-BFGS optimizer with clean data (noise = $0\%$), the $\mathbb{L}^{2}$ relative error of the variable coefficients $\alpha(t)$ and $\beta(t)$ decreased from $0.40\%$ to $0.28\%$ and $0.65\%$ to $0.45\%$, respectively. When we add noise levels of $5\%$, $10\%$, and $15\%$ into the training observations, the $\mathbb{L}^{2}$ relative error of the dispersion coefficient $\alpha(t)$ changes from $3\%, 0.550$$5\%$, and $0.57\%$ to $0.33\%$, $0.27\%$, and $0.33\%$, respectively, when we first use iterations of the Adam optimizer and then continue the optimization process with the L-BFGS optimizer. This scenario holds for the nonlinearity coefficient $\beta(t)$, where the $\mathbb{L}^{2}$ relative error changes from $0.61\%$, $0.67\%$, and $0.72\%$ to $0.42\%$, $0.40\%$, and $0.60\%$, when we add noise levels of $5\%$, $10\%$, and $15\%$ into the training observations, respectively. The predictions of the dispersion and nonlinearity coefficients $\alpha(t)$ and $\beta(t)$ under different noise levels are displayed in the left and middle panels of Fig. 18. One can see that our enhanced trained PINN model can accurately predict the dynamic behavior of both variable coefficients, even in the presence of noise intensities. The $\mathbb{L}^{2}$ relative errors of both the dispersion and nonlinearity variable coefficients $\alpha(t)$ and $\beta(t)$ for different noise intensities are illustrated in the right panel of Fig. 18, which shows that increasing the noise levels does not significantly impact prediction accuracy. The obtained results show that our enhanced PINN model can accurately predict the behavior of the variable coefficients $\alpha(t)$ and $\beta(t)$ in the system of model equations (1) with a small amount of training data.
@@
Figure 18: The predictions of the dispersion coefficient $\alpha(t)$ (left panel) and the nonlinearity coefficient $\beta(t)$ (middle panel), and the related prediction errors (right panel) under different noise intensities.
@@ 4 Conclusions
In this study, we proposed an enhanced PINN framework for data-driven soliton solutions and parameter discovery to the (2+ 1)-dimensional VC-CNLSE. PINN algorithms enhance the capabilities of neural networks by incorporating the physical knowledge, described by differential equations, into the training phase. Incorporating physical knowledge into the training phase allows the network to learn from both data and the underlying principles governing the problem, leading to more accurate and physically consistent predictions. Our proposed framework uses a locally adaptive activation function to accelerate the convergence speed of the training process for the underlying PINN model. By leveraging the residual-based adaptive refinement (RAR) strategy, we introduced a region-specific weighted loss function, demonstrating that different distributions of physical knowledge are required when training a PINN model for the (2+ 1)-dimensional VC-CNLSE. The minimization problems are carried out using iterations of the Adam optimizer followed by iterations of the L-BFGS optimizer. The efficiency and accuracy of the proposed framework are evaluated in capturing the dynamics of vector dark and anti-dark one- and two-soliton structures, using various soliton solutions and data-driven predictions. The density plots of the exact and predicted solutions are displayed, and the related evolution dynamics are illustrated through three-dimensional plots. We also compared the exact and predicted solutions at different temporal levels to show the accuracy of the trained model’s prediction. We performed data-driven parameter discovery for the (2+ 1)-dimensional VC-CNLSE, categorizing it into two classes: constant parameter discovery and variable parameter discovery. For constant parameter discovery, we employed an enhanced PINN model equipped with a single neural network to estimate the value of the constant parameter in the system of model equations. For variable parameter discovery, we used a dual-network strategy within the proposed PINN framework to capture the dynamic behavior of the variable dispersion and nonlinearity coefficients. As these coefficients are nonlinear functions of the temporal variable, this shows that our enhanced PINN framework is capable of capturing complex nonlinear variable coefficients. This robustness highlights the framework’s potential for broader applications in complex optical fiber systems. The obtained results confirm that the proposed enhanced PINN framework offers a powerful and versatile approach for solving high-dimensional and complex solitonic dynamics in optical communication systems. Its robustness, accuracy, and adaptability make it a promising tool for advancing research and applications in nonlinear optics and related fields.
Declaration of competing interest
The authors declare that there is no conflict of interest.
Data availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.